OpenStudy (anonymous):

NEED HELP! Factor this binomial completely: x^9 -y^9

OpenStudy (idku):

\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

You need to factor by making a substitution to create a more familiar form of the difference of two CUBES. Let z = x^3 so that you expression becomes z^3-z^3 ( because (x^3)^3 = z^3). Do you know the formula for factoring a^3-b^3?